# ToDo (#1): #2

### v2

@@ 3.8.1 Isochronization
The most likely deployment scenario is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea. Fig. 4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator conveyance in a new 600 meter vertical shaft. Unfortunately, the elevator is suitable only for personnel access, all heavy and bulky equipment must be brought down through the ramp. Fig. 4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter). Site-specific questions will be addressed in Volume III of this PDR and the following section aims at a general description of deployment strategies in an underground setting, pertaining only to the cyclotron. In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at other installation sites.
Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure.
@@ 4.1 Transport of Components
While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf. Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions. Improved scans must be performed as part of final installation planning, and will be included in volume III which will be site-specific to Yemilab.
As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site. The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig. 4.3 for bringing the magnet and vacuum liner pieces from the surface. These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [45].
Figure 4.3: A typical SPMT vehicle, extremely versatile and capable of extremely heavy loads over irregular terrain
@@
In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections. This has already been studied in Chapter 3 of the CDR document [46, 2]. The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site. The biggest challenge would be the ventilation scheme for the potting process. There is precedence for sectioned coils. For the large TRIUMF cyclotron [47], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways. The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
@@ 4.2 Assembly underground
The cyclotron will be assembled in a hall measuring 12 meters wide, 17 meters deep and 10 meters high. This hall was excavated to our specifications at the 90 degree bend in the access ramp to the LSC, as seen in Fig. 4.2. A flat area, roughly 10 m x 5 m at the corner will be the staging area for components transported from the surface. A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point. Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A. [48] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig. 4.6. It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons
Figure 4.7: Cyclotron and crane on cavern floor.
@@
Figure 4.9: Cycloton open and crane with cavern mesh - cut-view.
@@ 5.1 Summary
In this version 0.9 of the document, we presented the preliminary design of the IsoDAR compact cyclotron with RFQ injection, including ion source and low energy beam transport.
@@
The presented design study shows the feasibility of producing a high current (10 mA of protons through stripping of a 5 mA $\mathrm{H}_{2}^{+}$ beam) in a compact cyclotron. It further shows a convincing strategy to segment the cyclotron yoke for transportation into the underground environment to be assembled in situ at the final location.
@@
RF calculations show manageable power and cooling, the required radial voltage distribution in the accelerating gaps, and adequate room for vacuum getter pumps.
In a future update to this PDR document, we plan to show a full start-to-end beam dynamics study, based on the presented individual components’ beam simulations. We will also add a risk assessment, and additional details on the extraction system, based on the cited publications and ongoing design work. Finally, we will include a vacuum calculation based on MolFlow.
@@ 5.2 Acknowledgements
This work was primarily supported by NSF grant PHY-2012897. Preliminary work leading up to this was supported by NSF grants PHY-1912764 and PHY-1626069. Moon was supported by DOE grant DE-SC0024138. Winklehner was supported by the Heising-Simons Foundation.

### v3

@@ 3.9.6 Risks and Mitigation
The most likely deployment location for IsoDAR is at the Yemilab underground laboratory, 1 km below the surface adjacent to the Handuk iron mine in the Republic of Korea. Fig. 4.1 schematically shows Yemilab, accessed either by a 6 km meter ramp descending 750 meters at a 15% slope from a surface access point into the mine, or by an elevator which descends down a 600 meter vertical shaft. Unfortunately, the elevator is suitable only for personnel access. All heavy and bulky equipment must be brought down via the ramp. Fig. 4.2 details the area where the IsoDAR experiment would be installed, with the target only a few meters from the planned 3 kiloton LSC (Liquid Scintillator Counter). The following section provides a general description of deployment strategies in an underground setting, pertaining only to the cyclotron. In this document we use Yemilab as the example, noting that most of the issues discussed will arise in an equivalent manner at any alternative installation sites. A similar approach will be taken for the target system in Volume II of this PDR.
Figure 4.1: (a) Schematic of Handuk/Yemilab area, (b) Low resolution laser scan of Handuk mine and Yemilab access ramps, green area denotes high resolution scans of IsoDAR deployment area, shown in more detail in the next figure. From Ref. [kimYemilabNewUnderground2024].
@@ 4.1 Transport of Components
While high-resolution laser scans of the IsoDAR and LSC areas have been performed, (cf. Figures 4.8 and 4.9), the resolution of the laser scans of the access ramp areas are not of sufficient resolution to provide enough detailed information about the location or size of the possible constrictions. Improved scans must be performed as part of final installation planning in the Technical Design Report phase (TDR).
As seen in the previous chapter, the iron pieces that make up the cyclotron magnet, although heavy, are not in themselves large, so the pinch points are not likely to be a problem for bringing the magnet steel to the installation site. The plan is to use an SPMT vehicle (Self-Propelled Modular Transporter) shown in Fig. 4.3 for bringing the magnet and vacuum liner pieces from the surface. These electrically-powered vehicles are flexible in their size, the sets of 4 tires are independently steerable, so the vehicle is extremely maneuverable, and can easily accommodate the heaviest piece of magnet steel [SPMT].
Figure 4.3: A typical SPMT vehicle, a versatile way to transport heavy loads over irregular terrain.
@@
In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections. This has already been studied in Chapter 3 of the CDR document [alonso_isodaryemilab_2021, alonso_isodaryemilab_2022]. The heavy winding machine and fixtures would need to be transported, but could be broken down into parts that will fit through the ramps, then assembled on site. The biggest challenge would be the ventilation scheme for the potting process. There is precedence for sectioned coils. For the large TRIUMF cyclotron [TRIUMF], coils are separated into six different segments as the outer diameter of the coils is almost 20 meters, making the transport of fully assembled coils essentially impossible over conventional roadways. The inconvenience for a segmented coil, is that the ends of every turn of every section must be spliced, substantially increasing the cost and decreasing the efficiency of operation, as the coil must have as few turns as possible with very high current flowing through each turn.
@@ 4.2 Assembly underground
The cyclotron will be assembled in a hall measuring 12 meters wide, 17 meters deep and 10 meters high. This hall was excavated to our specifications at the 90 degree bend in the access ramp to the LSC, as seen in Fig. 4.2. A flat area, roughly 10 m x 5 m at the corner will be the staging area for components transported from the surface. A bridge crane riding on tracks will be used to rig parts to the cyclotron assembly point. Conversations with a crane manufacturer, Industrias Electromecánicas GH, S.A. [GHcranes] gave us assurances that a 40-ton capacity crane, riding on rails installed on the cavern floor would be quite practical (see Fig. 4.6. It could be configured to fit the dimensions of the cavern, and broken into modular parts that could be transported to the site.
Figure 4.6: Examples of GH gantry cranes, capacities: foreground 20 tons, background 260 tons. From Ref. [GHcranes].
@@
Figure 4.9: Cyclotron open and crane with cavern mesh - cut-view.
4.2.1 Risks and Mitigations
Risk: Coil Size. The cyclotron coil will likely not fit through all of the sections of the main access ramp at Yemilab in horizontal orientation. A special mounting fixture capable of holding the coil packages at an angle should be able to bring these coils through diagonally. However, there is some uncertainty about possible choke points in the main access ramp to the Yemilab facility.
Mitigation: In the unlikely case the individual pancakes do not fit through the constriction points, the alternative is to wind and pot the coils underground, or make them in sections, as discussed above. Both come with drawbacks or additional cost, but are manageable.
@@ 5.1 Summary
In this version 2.0 of the document, we presented the preliminary design of the IsoDAR compact cyclotron with RFQ direct injection, including ion source, low energy beam transport, electrostatic extraction channels and some thoughts on typical magnetic extraction channels. This version presents significant updates over the previously published CDR and also version 0.9 of this PDR. Most notably, we included first information about the upgraded MIST-2 ion source, a much optimized spiral inflector, a realistic cyclotron magnet and RF cavity design and particle-in-cell simulations thereof. Over version 0.9 of this document, we added detailed subsections on the simulations and subsections on risks associated with the different components as well as mitigation strategies.
@@
The presented design study shows the feasibility of producing a high CW $\mathrm{H}_{2}^{+}$ beam current (10 mA of protons through stripping of a 5 mA $\mathrm{H}_{2}^{+}$ beam) in a compact cyclotron. It further shows a convincing strategy to segment the cyclotron yoke for transportation into the underground environment to be assembled in situ at the final location. Below, we summarize our designs and findings for the different subsystems (ion source and LEBT, RFQ, cyclotron, underground installation).
5.1.1 Ion Source and LEBT Summary
Based on literature and in-depth simulations using well-established codes, we have determined that a filament-driven multicusp ion source presents an attractive solution to producing 10-12 mA of DC $\mathrm{H}_{2}^{+}$ beam current with high purity and low emittance. The filament lifetime is the main drawback. We presented our design and measurements with such an ion source using our own MIST-1 prototype and first tests of the upgraded MIST-2 components. Our measurements agree extremely well with simulations and, although we have only reached about half of the design current with the MIST-1 prototype, we are confident that we have addressed all the “lessons learned” from it in our MIST-2 design and foresee no showstoppers in achieving the needed beam current with it. We estimate the filament lifetime to be about a week of continuous running, which would let us achieve the required 90% up-time. As a fallback, flat-field ECR ion sources are a tried-and-tested alternative, albeit with the (manageable) caveats of co-producing significant amounts of protons and having a higher emittance.
The LEBT is a straightforward, highly compact design that includes safety mechanisms, beam positioning, and beam diagnostics. Our simulations include robustness tests and failure-mode simulations showing its capability to handle these. The beam is well matched into the RFQ. Parts of the LEBT have been manufactured with the remainder to be delivered at the same time as the RFQ. Experimental verification is thus forthcoming.
5.1.2 RFQ Summary
Here and through references to published papers, we presented the technical design of a split-coaxial RFQ of 28 cm diameter, operating at a frequency of 32.8 MHz (matched to the cyclotron frequency). The design includes thermal and mechanical vibration studies and full 3D PIC beam dynamics calculations including all fringe fields, using the LEBT output as input beam.
The RFQ is currently being manufactured by Bevatech, GmbH in Germany. An experimental campaign has been started to put together ion source, LEBT and RFQ, and ultimately a 1.5 MeV/amu test cyclotron (the HCHC-1.5) for the ultimate demonstration of $\mathrm{H}_{2}^{+}$ acceleration, RFQ direct injection, and establishment of vortex motion.
5.1.3 Cyclotron Design Summary
We presented a realistic cyclotron mechanical and electrical design consisting of the main magnet, the RF Cavities, and vacuum chamber. The magnetic and electric fields calculated from these models by FEM and BEM were used to perform highly realistic PIC simulations using the OPAL code.
Simulations merging the RFQ output with the spiral inflector in the central region have now also been performed; completing a simulation chain running seamlessly from initial plasma generation in the ion source into the central region of the cyclotron where acceleration will begin. Simulations of the acceleration beginning with the end of the central region and concluding with extraction at the target energy of 60 MeV/amu have been performed using an input beam that was informed by a standalone central region design. Connecting the latest spiral inflector simulations with a modified version of this central region inside the new magnet model will yield a full start-to-end simulation of the IsoDAR experiment.
This section also discussed several improvements that have been explored to increase the efficiency, safety, and beam quality that emerges from the inflector into the cyclotron.
@@ 5.1.3 Cyclotron Design Summary
Calculations performed with the full 3D model of the RF cavity showed manageable power and cooling. The required radial voltage distribution in the accelerating gaps, and adequate room for vacuum getter pumps.
5.1.4 Installation Summary
We described the layout of the Yemilab facility and how to transport the cyclotron steel components using SPMTs. In the cyclotron cavern, assembly will be aided by a gantry crane that can be disassembled after cyclotron installation to make room for the peripheral devices.
We identified the main risk as the cyclotron coil size being on the edge of what can be transported through the main access ramp at Yemilab. A careful investigation of possible choke points is needed for the TDR. Possible mitigation is to wind the coil in place or to adopt a split coil design as we described in the CDR.
For the next level of detail, we are working hand in hand with colleagues from Yemilab to identify further issues, devise an installation and removal plan, as well as overall integration plans with the facility (power, HVAC, controls, etc.).
5.2 Outlook
A future Technical Design Report will have to include a number of additional details, among them a full start-to-end simulation using as few beam dynamics codes as possible, benchmarked against a second set of codes; robustness analysis; a vacuum calculation (most likely using the CERN-developed code MolFlow); specifics of assembly and removal as well as shielding of the cyclotron cavern; RF system details; a commissioning strategy; data acquisition for beam diagnostics; and the cyclotron control system.
We have investigated a number of these topics and published results thereon, but including them here would have gone beyond the level of a PDR. Others are current and ongoing studies and publications are forthcoming (well before the final TDR, much like we did here).
Most notably, we are in the process of putting together the RFQ-Direct Injection Project (RFQ-DIP), combining the MIST-2 ion source (currently commissioning), the RFQ (currently being machined), and the HCHC-1.5 cyclotron (parts being ordered), as demonstrator of all novel concepts in the HCHC-60 cyclotron design.
@@ 5.3 Acknowledgements
This work was primarily supported by NSF grant PHY-2012897. Preliminary work leading up to this was supported by NSF grants PHY-1912764 and PHY-1626069. Work for V1.0 was supported by DOE grants DE-SC0024138 and DE-SC0024914 and NSF grant PHY-2411745. Winklehner was also supported by the Heising-Simons Foundation. Winkler thankfully acknowledges the support from the MIT UROP office.
[…]